# In-N-Out: A Parameter-Level API Graph Dataset for Tool Agents

### v2

@@ Abstract
Tool agents—LLM-based systems that interact with external APIs—offer a way to execute real-world tasks. However, as tasks become increasingly complex, these agents struggle to identify and call the correct APIs in the proper order. To tackle this problem, we investigate converting API documentation into a structured API graph that captures API dependencies and leveraging it for multi-tool queries that require compositional API calls. To support this, we introduce In-N-Out, the first expert-annotated dataset of API graphs built from two real-world API benchmarks and their documentation. Using In-N-Out significantly improves performance on both tool retrieval and multi-tool query generation, nearly doubling that of LLMs using documentation alone. Moreover, graphs generated by models fine-tuned on In-N-Out close 90% of this gap, showing that our dataset helps models learn to comprehend API documentation and parameter relationships. Our findings highlight the promise of using explicit API graphs for tool agents and the utility of In-N-Out as a valuable resource. We will release the dataset and code publicly.
@@ 1 Introduction
While prior studies have explored training tool agents on large datasets of multi-tool queries Qin et al. (2024); Shim et al. (2025), a more natural and general solution, akin to what human developers do, is to use API documentation to identify API dependencies. However, since LLMs struggle to comprehend parameter-level dependencies from noisy, real-world API documentation (§4), we focus on a relatively underexplored approach: converting API documentation into an explicit API graph, enabling LLMs to use this structured information. Some prior studies that have attempted to use API graphs have focused primarily on limited domains Wang et al. (2024) or automatic graph construction with heuristics Liu et al. (2024b); Shen et al. (2024), leaving the true potential of accurate API graphs in real-world scenarios unexplored. Since manual graph construction is not scalable for arbitrary API sets, a high-quality dataset is needed to train and evaluate a documentation-to-graph conversion module and ensure its generalizability to unseen APIs. Our approach is arguably more generalizable and scalable than simply training a tool agent on large multi-tool tasks, and the graphs can even facilitate the generation of multi-tool queries by providing the underlying API composition for complex queries (Figure 1b).
Against this backdrop, we introduce In-N-Out, an API graph dataset constructed through expert annotation. In-N-Out represents both APIs and their parameters as nodes, with directed edges indicating when an output from one API can serve as a valid input to another. The APIs are sourced from AppWorld Trivedi et al. (2024) and NESTful Basu et al. (2025), spanning 550 APIs, and over 30,000 parameter-level edges, thus covering diverse and realistic tool-use scenarios.
@@ 2.1 Tool Retrieval and Planning
As real-world tasks increasingly require multiple APIs in sequence, both accurate retrieval and correct input provision become essential (Patil et al., 2025; Xu et al., 2024; Shi et al., 2025; Du et al., 2024; Lumer et al., 2024). To support these capabilities, previous work has explored several approaches: refining API documentation to make it more interpretable for LLMs (Hsieh et al., 2023; Yuan et al., 2025; Chen et al., 2024), fine-tuning LLMs on query-API call pairs (Qin et al., 2024; Patil et al., 2024), and learning tool-specific embeddings (Hao et al., 2023; Wang et al., 2025). However, these strategies often fail to capture the underlying dependencies among APIs, while also requiring high training costs and generalizing poorly to unseen APIs.
To address this, recent studies have explored graph-based approaches that represent API relationships explicitly to support tool retrieval (Liu et al., 2024b, a; Chen et al., 2025). GraphRAG-Tool Fusion (Lumer et al., 2025) constructs graphs by modeling both API- and parameter-level dependencies, but relies on synthetic APIs that lack the complexity of real-world environments. SoAy (Wang et al., 2024) uses real APIs but is limited to academic information-seeking tasks with only seven APIs and simple parameter structures, falling short of the real-world requirement to combine many APIs across diverse domains. In contrast, we construct expert-verified graphs over real APIs across 25 domains, capturing richer relationships and enabling reliable tool retrieval in real-world settings.
@@ 2.2 Multi-Tool Query Generation
To overcome these issues, recent work has leveraged graph structures to systematically generate multi-tool queries Wang et al. (2024). TaskBench Shen et al. (2024) constructs graphs by matching parameter data types (e.g., image, text), and samples subgraphs to generate queries. ToolDial Shim et al. (2025) connects APIs through keyword matching and embedding similarity, and generates queries by traversing pairs of connected nodes. While scalable, these approaches often connect semantically incompatible APIs or capture only coarse relations. In contrast, our work constructs parameter-level API graphs through expert annotation, providing accurate ground-truth connections that support valid query generation reflecting genuine dependencies. Moreover, models trained on these graphs can generalize to unseen APIs, enabling scalable inference of new dependencies and corresponding query generation.
@@ 3 In-N-Out Dataset
Our annotations are built on two realistic and diverse datasets: NESTful(v1) Basu et al. (2025) and AppWorld Trivedi et al. (2024). NESTful(v1) contains 38 real-world APIs sourced from RapidAPI Hub [^1] https://rapidapi.com/hub across 15 domains (e.g., weather, music, news), along with 85 multi-tool queries. We extend it with 55 additional APIs from the same domains to enrich the graph with more diverse relationships. AppWorld offers a simulated environment for 9 real-world apps (e.g., amazon, venmo, spotify) with 457 APIs and 750 multi-tool tasks, providing a rich benchmark for compositional API usage. Based on these two datasets, we construct graphs that serve as a valuable resource for effective API composition.

### v3

@@ Abstract
Tool agents—LLM-based systems that interact with external APIs—offer a way to execute real-world tasks. However, as tasks become increasingly complex, these agents struggle to identify and call the correct APIs in the proper order. To tackle this problem, we investigate converting API documentation into a structured API graph that captures API dependencies and leveraging it for multi-tool queries that require compositional API calls. To support this, we introduce In-N-Out, the first expert-annotated dataset of API graphs built from two real-world API benchmarks and their documentation. Using In-N-Out significantly improves performance on both tool retrieval and multi-tool query generation, nearly doubling that of LLMs using documentation alone. Moreover, graphs generated by models fine-tuned on In-N-Out close 90% of this gap, showing that our dataset helps models learn to comprehend API documentation and parameter relationships. Our findings highlight the promise of using explicit API graphs for tool agents and the utility of In-N-Out as a valuable resource. We release our dataset and code at https://github.com/holi-lab/In-N-Out-API-Graph.
@@ 1 Introduction
While prior studies have explored training tool agents on large datasets of multi-tool queries Qin et al. (2024); Shim et al. (2025), a more natural and general solution, akin to what human developers do, is to use API documentation to identify API dependencies. However, since LLMs struggle to comprehend parameter-level dependencies from noisy, real-world API documentation (§4), we focus on a relatively underexplored approach: converting API documentation into an explicit API graph, enabling LLMs to use this structured information. Some prior studies that have attempted to use API graphs have focused primarily on limited domains Wang et al. (2025b) or automatic graph construction with heuristics Liu et al. (2024b); Shen et al. (2024), leaving the true potential of accurate API graphs in real-world scenarios unexplored. Since manual graph construction is not scalable for arbitrary API sets, a high-quality dataset is needed to train and evaluate a documentation-to-graph conversion module and ensure its generalizability to unseen APIs. Our approach is arguably more generalizable and scalable than simply training a tool agent on large multi-tool tasks, and the graphs can even facilitate the generation of multi-tool queries by providing the underlying API composition for complex queries (Figure 1b).
Against this backdrop, we introduce In-N-Out, an API graph dataset constructed through expert annotation. In-N-Out represents both APIs and their parameters as nodes, with directed edges indicating when an output from one API can serve as a valid input to another. The APIs are sourced from AppWorld Trivedi et al. (2024) and NESTful(v1) Basu et al. (2025), spanning 550 APIs, and over 30,000 parameter-level edges, thus covering diverse and realistic tool-use scenarios.
@@ 2.1 Tool Retrieval and Planning
As real-world tasks increasingly require multiple APIs in sequence, both accurate retrieval and correct input provision become essential (Patil et al., 2025; Xu et al., 2024; Shi et al., 2025; Du et al., 2024; Lumer et al., 2024). To support these capabilities, previous work has explored several approaches: refining API documentation to make it more interpretable for LLMs (Hsieh et al., 2023; Yuan et al., 2025; Chen et al., 2024), fine-tuning LLMs on query-API call pairs (Qin et al., 2024; Patil et al., 2024), and learning tool-specific embeddings (Hao et al., 2023; Wang et al., 2025a). However, these strategies often fail to capture the underlying dependencies among APIs, while also requiring high training costs and generalizing poorly to unseen APIs.
To address this, recent studies have explored graph-based approaches that represent API relationships explicitly to support tool retrieval (Liu et al., 2024b, a; Chen et al., 2025). GraphRAG-Tool Fusion (Lumer et al., 2025) constructs graphs by modeling both API- and parameter-level dependencies, but relies on synthetic APIs that lack the complexity of real-world environments. SoAy (Wang et al., 2025b) uses real APIs but is limited to academic information-seeking tasks with only seven APIs and simple parameter structures, falling short of the real-world requirement to combine many APIs across diverse domains. In contrast, we construct expert-verified graphs over real APIs across 25 domains, capturing richer relationships and enabling reliable tool retrieval in real-world settings.
@@ 2.2 Multi-Tool Query Generation
To overcome these issues, recent work has leveraged graph structures to systematically generate multi-tool queries Wang et al. (2025b). TaskBench Shen et al. (2024) constructs graphs by matching parameter data types (e.g., image, text), and samples subgraphs to generate queries. ToolDial Shim et al. (2025) connects APIs through keyword matching and embedding similarity, and generates queries by traversing pairs of connected nodes. While scalable, these approaches often connect semantically incompatible APIs or capture only coarse relations. In contrast, our work constructs parameter-level API graphs through expert annotation, providing accurate ground-truth connections that support valid query generation reflecting genuine dependencies. Moreover, models trained on these graphs can generalize to unseen APIs, enabling scalable inference of new dependencies and corresponding query generation.
@@ 3 In-N-Out Dataset
Our annotations are built on two realistic and diverse datasets: NESTful(v1) Basu et al. (2025) and AppWorld Trivedi et al. (2024). NESTful contains 38 real-world APIs sourced from RapidAPI Hub [^1] https://rapidapi.com/hub across 15 domains (e.g., weather, music, news), along with 85 multi-tool queries. We extend it with 55 additional APIs from the same domains to enrich the graph with more diverse relationships. AppWorld offers a simulated environment for 9 real-world apps (e.g., amazon, venmo, spotify) with 457 APIs and 750 multi-tool tasks, providing a rich benchmark for compositional API usage. Based on these two datasets, we construct graphs that serve as a valuable resource for effective API composition.
@@ 6 Limitations
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant (RS-2024-00333484) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) under the Leading Generative AI Human Resources Development grant (IITP-2026-RS-2024-00397085) and the grant (RS-2025-02215122, Development and Demonstration of Lightweight AI Model for Smart Homes), all of which are funded by the Korean government (MSIT). This work was also supported by the National Supercomputing Center with supercomputing resources including technical support (KSC 2024-CRE-0554, KSC-2025-CRE-0332).